OpenStudy (anonymous):

what is the inverse function of f(x)= x+6?

OpenStudy (zzr0ck3r):

y = x+6 replace x with y and y with x x = y+6 solve for y y = x-6

OpenStudy (zzr0ck3r):

@stephanied2013 ok?

OpenStudy (anonymous):

got it , thanks !